OpenStudy (anonymous):

I have to complete statements. I have 4 to do but here is one. A(1) = 2 = 2 * 3^ (it has an exponent but I think I need to find the exponent somehow?) Could you explain to me how to complete this? I have four more.

OpenStudy (anonymous):

@AravindG :)

OpenStudy (aravindg):

sorry I am late ok this one is easy ..call the unknown variable as x 2=2*(3^x) 2=2 so 3^x should be 1 3^x=1 x=0 :)

OpenStudy (anonymous):

thank you so much!!

OpenStudy (aravindg):

yw :) click best response if that helped :)